OpenStudy (anonymous):

Compute Z=(1+i)(1+2i)(1+3i).

OpenStudy (anonymous):

(1+i)(1+2i) = 1 + 2i + i +2(-1) = -1 + 3i (-1 + 3i)(1 + 3i) = -1 -3i + 3i - 9 = -10

OpenStudy (anonymous):

because i^2 = -1 by definition